OpenStudy (anonymous):

Given the function f(x) = The quantity of 3x minus 4, divided by 5, which of the below expressions is correct? f−1(x) = The quantity of 5x plus 4, divided by 3. f−1(x) = The quantity of 5x minus 4, divided by 3. f−1(x) = The quantity of negative 3x minus 4, divided by 5. f−1(x) = The quantity of 4 minus 3x, divided by 5.

OpenStudy (lynfran):

solve for x y=(3x-4)/5 5y=3x-4 5y+4=3x (5y+4)/3=x switch the x" and y'' (5x+4)/3=y f(x)=(5x+4)/3

OpenStudy (lynfran):

f'(x)=(5x+4)/3